# Continual Multi-Robot Learning from Black-Box Visual Place Recognition Models

### v1

@@ IV-A Experimental Setup
    Sequential teacher interaction (up to 3 teachers)
@@
    27 NCLT dataset sessions
    • 1 test session: “2012/08/04”
    • 1 training session for embedding model: “2012/04/29”
    • 25 sessions for student/teacher VPR model training
@@ 27 NCLT dataset sessions
    Teacher/Student VPR models
    • MLP with a hidden layer of 4,096 dimensions
@@
    6 distinct knowledge transfer scenarios
@@
      • After the student’s supervised learning
      • After knowledge transfer from the 1st teacher
      • After knowledge transfer from the 2nd teacher
      • After knowledge transfer from the 3rd teacher
@@
      • The number of samples $N$ used for KT
    • Knowledge retention
      • Avoidance of catastrophic forgetting
    • Computational efficiency
@@
Figure 3: Example of experimental setup. In each scenario, at stage $i=0$, the student robot trains the VPR model via supervised learning using the data from the place classes it has experienced (blue boxes). In subsequent stages $i=1,2,3$, every time the student encounters a new teacher, it retrains the VPR model via knowledge transfer using the data of the place classes the teacher has experienced (yellow boxes).
@@
Several detailed setup configurations are shown below: (1) The student robot and teacher robots have experienced $K$ place classes and have trained a place classifier in a supervised manner using the full training set for these experienced place classes from the corresponding session. Unless otherwise specified, following the standard continual learning protocol, once the training is completed, the training set is not retained and is discarded for all BB-MIA strategies except Replay, Prior, and Mixup. (2) Session IDs from 0 to 24 were assigned to 25 different sessions, in order of the navigation dates, starting from the earliest. (3) To evaluate performance across different student-teacher combinations, we introduce different scenarios for $j=0,1,\dots,5$. In the $j$-th scenario, the student and the three teachers are distinguished by model ID $i$. Specifically, $i=0$ represents the student model, while $i=1,2,3$ correspond to the $i$-th teacher. Each $i$-th model experiences $K$ random place classes during the $((6i+j)\mod 25)$-th session among the 25 sessions. By default, $K=10$. (4) For simplicity, the number of samples appearing in this section will refer to the number of samples per place class, rather than the total number across all place classes. (5) When the student encounters a teacher, (pseudo) training samples are reconstructed from the teacher model for place classes known to the teacher, while for place classes known only to the student, training samples are reconstructed from the student model. (6) In the Replay, Prior and Mixup strategy, the number of mixed replay samples $R$ per class was set to $R=1$ by default. (7) RRF vectors were approximated using a sparse $k$-hot RRF ($k=10$) following the method in [19].
@@
The computational cost of this experiment was lightweight, with training times of tens of seconds for MLP models and about 25 milliseconds per sample for question generation. Additionally, the KT cost of transferring a 100-dimensional $k$-hot RRF sample was less than 128 bits.
IV-B Results and Discussions
First, we evaluated the basic performance of the proposed method. We examined the performance at the initial stage when the student was trained with supervised learning, as well as at later stages when the student continued learning through knowledge transfer from new teachers.
@@ IV-B Results and Discussions
At the initial stage, when the student had not yet encountered any teachers, its performance was predictably poor. This was because test sessions included questions from unseen place classes, making it evident that even the most capable student would struggle to provide accurate answers for such test samples.
Figure 4: Top-1 accuracy vs. KT cost ($N$).
@@
Replay Strategy: When the number of samples was sufficiently large, this strategy remained unaffected by catastrophic forgetting [22] and achieved the highest performance across all experiments. However, it comes at the cost of not adhering to the principles of CL. The Prior strategy, a variant of the Replay strategy, which queries the teacher using the student’s training samples, did not perform well. This may be due to the low probability of the student’s training sample coinciding with the teacher’s training sample by chance.
@@
Entropy Strategy: While being still simple, it performed comparably or better than the RR strategy, especially excelling at generating elite samples when $N$ was small.
Mixup Strategy: This strategy balances generalization and cost-effectiveness.
As mentioned earlier, both Replay and Mixup do not satisfy the assumption of a black-box teacher and are used solely for benchmarking purposes. Although Mixup requires retaining a small number ($R$) of training samples, its knowledge transfer cost is significantly lower than that of Replay, mitigating catastrophic forgetting while achieving performance close to that of the Replay strategy.
V Conclusion and Future Work
We formulated the problem of continual multi-robot learning for visual place recognition (VPR) as teacher-student knowledge transfer and proposed exploiting membership inference attacks (MIA), the only major privacy attack capable of handling black-box teachers. To enable the use of MIA for knowledge transfer, we focused on the key challenge of its low sampling efficiency, which acts as a bottleneck. We presented a method that leverages the student model’s prior knowledge to achieve practical sampling efficiency. Through extensive experiments, we investigated the relationship between VPR performance, sampling efficiency, and computational efficiency, demonstrating the significant effectiveness of our proposed approach.
In this study, we focused on a fundamental investigation, addressing a near-worst-case scenario where all teachers are black boxes and all robots have low performance. For practical deployment, future research must urgently explore heterogeneous multi-robot systems that include white-box teachers and robots equipped with specialized continual learning capabilities. Furthermore, while this study adopted a highly naive grid-based partitioning for place definition, prior research suggests that a straightforward extension using more advanced place definition methodologies could dramatically improve visual place recognition performance. Another promising research direction is extending beyond single-image-based visual place recognition to robust self-localization using image sequences and particle filtering, which offers a guaranteed performance boost.

### v2

@@ V-A Experimental Setup
    Sequential teacher interaction (up to 3 teachers).
@@
    27 NCLT dataset sessions:
    • 1 test session: “2012/08/04”;
    • 1 training session for embedding model: “2012/04/29”;
    • 25 sessions for student/teacher VPR model training.
@@ 27 NCLT dataset sessions:
    Teacher/Student VPR models:
    • MLP with a hidden layer of 4,096 dimensions.
@@
    6 distinct knowledge transfer scenarios.
@@
      • After the student’s supervised learning;
      • After knowledge transfer from the 1st teacher;
      • After knowledge transfer from the 2nd teacher;
      • After knowledge transfer from the 3rd teacher.
@@
      • The number of pseudo-samples $N$ used for communicative KT.
    • Knowledge retention:
      • Avoidance of catastrophic forgetting.
    • Computational efficiency.
@@
Figure 3: Example of experimental setup. In each scenario, at stage $i=0$, the student robot trains the VPR model via supervised learning using the data from the place classes it has experienced (blue boxes). In subsequent stages $i=1,2,3$, every time the student encounters a new teacher, it retrains the VPR model via communicative knowledge transfer using the data of the place classes the teacher has experienced (yellow boxes).
@@
Several detailed setup configurations are shown below: (1) The student robot and teacher robots have experienced $K$ place classes and have trained a place classifier in a supervised manner using the full training set for these experienced place classes from the corresponding session. Unless otherwise specified, following the standard continual learning protocol, once the training is completed, the training set is not retained and is discarded for all BB-MIA strategies except Replay, Prior, and Mixup.
(2) Session IDs from 0 to 24 were assigned to 25 different sessions, in order of the navigation dates, starting from the earliest.
(3) To evaluate performance across different student–teacher combinations, we introduce different scenarios for $j=0,1,\dots,5$. In the $j$-th scenario, the student and the three teachers are distinguished by model ID $i$. Specifically, $i=0$ represents the student model, while $i=1,2,3$ correspond to the $i$-th teacher. Each $i$-th model experiences $K$ random place classes during the $((6i+j)\bmod 25)$-th session among the 25 sessions. By default, $K=10$.
(4) For simplicity, the number of samples appearing in this section will refer to the number of samples per place class, rather than the total number across all place classes.
(5) When the student encounters a teacher, (pseudo) training samples are reconstructed from the teacher model for place classes known to the teacher, while for place classes known only to the student, training samples are reconstructed from the student model.
(6) In the Replay, Prior and Mixup strategy, the number of mixed replay samples $R$ per class was set to $R=1$ by default.
@@
(7) RRF vectors were approximated using a sparse $k$-hot RRF ($k=10$) following the method in [27]. The computational cost of this experiment was lightweight, with training times of tens of seconds for MLP models and about 25 milliseconds per sample for question generation. Additionally, the KT cost of transferring a 100-dimensional $k$-hot RRF sample was less than 128 bits.
V-B Results and Discussions
@@
First, we evaluated the basic performance of the proposed CCL-based method. We examined the performance at the initial stage when the student was trained with supervised learning, as well as at later stages when the student continued learning through communicative knowledge transfer from new teachers. At the initial stage, when the student had not yet encountered any teachers, its performance was predictably poor. This was because test sessions included questions from unseen place classes, making it evident that even the most capable student would struggle to provide accurate answers for such test samples.
Figure 4: Top-1 accuracy vs. communicative KT cost (the number of pseudo-samples NN).
@@
Replay Strategy: When the number of samples was sufficiently large, this strategy remained largely unaffected by catastrophic forgetting [16] and achieved the highest performance across all experiments. However, it comes at the cost of not adhering to the principles of strictly data-free CCL. The Prior strategy, a variant of the Replay strategy, which queries the teacher using the student’s training samples, did not perform well. This may be due to the low probability of the student’s training sample coinciding with the teacher’s training sample by chance.
@@
Entropy Strategy: While still simple, it performed comparably or better than the RR strategy, especially excelling at generating elite samples when $N$ was small.
Mixup Strategy: This strategy balances generalization and cost-effectiveness. As mentioned earlier, both Replay and Mixup do not satisfy the strict assumption of a black-box teacher and are used solely for benchmarking purposes. Although Mixup requires retaining a small number ($R$) of training samples, its knowledge transfer cost is significantly lower than that of Replay, mitigating catastrophic forgetting while achieving performance close to that of the Replay strategy.
VI Conclusion and Future Work
We formulated the problem of continual multi-robot learning for visual place recognition (VPR) as a communicative teacher–student knowledge transfer problem, and introduced Continual Communicative Learning (CCL), a data-free multi-robot framework in which a traveler robot acquires knowledge from black-box teacher VPR models via Membership Inference Attacks (MIA). To enable the use of MIA for knowledge transfer, we focused on the key challenge of its low sampling efficiency, which acts as a communication bottleneck. We presented a method that leverages the student model’s prior knowledge to achieve practical sampling efficiency in high-dimensional VPR settings. Through extensive experiments, we investigated the relationship between VPR performance, sampling efficiency, and computational efficiency, demonstrating the significant effectiveness of our CCL-based approach.
In this study, we focused on a fundamental investigation, addressing a near-worst-case scenario where all teachers are black boxes and all robots have relatively low performance. For practical deployment, future research must explore heterogeneous multi-robot systems that include white-box teachers and robots equipped with specialized continual learning capabilities [7, 8]. Furthermore, while this study adopted a naive grid-based partitioning for place definition, prior research suggests that a straightforward extension using more advanced place definition methodologies could dramatically improve visual place recognition performance [23, 3, 4]. Another promising research direction is extending beyond single-image-based VPR to robust self-localization using image sequences and particle filtering, which offers a guaranteed performance boost. Finally, CCL could be combined with broader communicative learning and multi-agent communication frameworks [9, 19, 20] to enable richer, protocol-aware knowledge exchange among large robot collectives.